The source of bitterness in dark chocolate is the compound theobromine, an alkaloid present in cocoa beans. In a sample of dark chocolate there are 6.52 x 10^20 molecules of theobromine present. How many milligrams of theobromine are present in the sample? The molecular formula for theobromine is C7H8N4O2.
1
Expert's answer
2019-10-15T08:35:56-0400
"n=\frac{N}{N_A}= \frac{6.52\times10^{20}}{6.022\times 10^{23}}=1.08\times10^{-3}mol"

"m=M\times n"

"M(C_7H_8N_4O2)=180.164 \frac{g}{mol}"



"m(C_7H_8N_4O_2)= 180.164\times1.08\times10^{-3}= 0.195 g =195 mg"